The reaction rate of nitrogen monoxide and oxygen into nitrogen dioxide is given by the
expression
r = k [NO]2 [O2],
where r = reaction rate
k = reaction rate coefficient, which increases with increasing temperature.
a) What is the increase in the reaction rate when the concentration of nitrogen oxide is
doubled and the concentration of oxygen is tripled?
b) How much do you need to increase the nitrogen concentration to affect a nine-fold
increase in reaction rate, without changing other factors tha
1
Expert's answer
2011-02-03T06:24:42-0500
https://upload.wikimedia.org/math/4/9/3/493ba8a5537c9100593d9360d9a59543.png

a) r = k [2*NO]2 [3*O2] = 4*3 k [NO]2 [O2] = 12 r - the reaction rate will increase in 12 times.
b) you have to increase the concentration of the nitrogen in 3 times: 9r = k[3*NO]2[O2] = 32 k [NO]2[O2]